# GOAL: A Generalist Combinatorial Optimization Agent Learning

### v2

@@ Abstract
Machine Learning-based heuristics have recently shown impressive performance in solving a variety of hard combinatorial optimization problems (COPs). However they generally rely on a separate neural model, specialized and trained for each single problem. Any variation of a problem requires adjustment of its model and re-training from scratch. In this paper, we propose GOAL (for Generalist combinatorial Optimization Agent Learning), a generalist model capable of efficiently solving multiple COPs and which can be fine-tuned to solve new COPs. GOAL consists of a single backbone plus light-weight problem-specific adapters for input and output processing. The backbone is based on a new form of mixed-attention blocks which allows to handle problems defined on graphs with arbitrary combinations of node, edge and instance-level features. Additionally, problems which involve heterogeneous types of nodes or edges are handled through a novel multi-type transformer architecture, where the attention blocks are duplicated to attend the meaningful combinations of types while relying on the same shared parameters. We train GOAL on a set of routing, scheduling and classic graph problems and show that it is only slightly inferior to the specialized baselines while being the first multi-task model that solves a wide range of COPs. Finally we showcase the strong transfer learning capacity of GOAL by fine-tuning it on several new problems. Our code is available at https://github.com/naver/goal-co/.
@@ 1 Introduction
Combinatorial Optimization (CO) problems are extensively studied in the Operations Research community, because they are both challenging in terms of theoretical complexity and central in many applications such as transportation, logistics and finance. Traditional CO methods include exact solvers (typically based on integer programming [korte_combinatorial_2018]) and approximation algorithms [vazirani_approximation_2003] which provide optimality or approximation guarantees, respectively, but do not scale; and heuristic algorithms [boussaid_survey_2013] which produce good-quality solutions faster but without guarantees. Effective heuristics heavily rely on expert knowledge and are generally problem-specific. A more recent and promising family of methods for CO is based on Machine Learning, either to learn the hyper-parameters of classic CO solvers, or to directly learn heuristics in an end-to-end fashion [bengio_machine_2021, cappart_combinatorial_2021]. In particular constructive neural heuristics have been successful at solving a variety of CO problems, including routing [bello_neural_2017, nazari_reinforcement_2018, kool_attention_2019, kwon_pomo_2020], scheduling [zhang_learning_2020, kwon_matrix_2022] and graph problems [khalil_learning_2017, qiu_dimes_2022]. However, in most cases, a separate model is specialized and trained for each problem, and sometimes even for each distribution of instances within a problem, especially for varying instance sizes. While much effort has been devoted to improving the generalization to different instance distributions [manchanda_generalization_2022, zhou_towards_2023, son_meta-sage_2023, drakulic_bq-nco_2023], generalization to different problems is less studied. Indeed, although there are strong neural heuristics for many CO problems, any variation of these problems requires adjustment of their model and re-training from scratch. While it is true that some problem-specific model tailoring may be required, since the structure of problem instances and solutions are very diverse, many problems are related and it would be interesting to exploit their commonalities. In this paper, we seek to answer the following question: Can we design and train a single generalist model so that it can be efficiently adapted to solve a variety of CO problems? The idea of pretraining a unified model on some pretext task so that it can be efficiently fine-tuned to several downstream tasks has been extremely successful in the language and vision domains [devlin_bert_2019, brown_language_2020, radford_learning_2021]. In these domains, a self-supervised pretraining task exploits the intrinsic structure and regularities of images and text to learn generic representations. In contrast, in CO, while most problem instances can be modeled as graphs, there is no generic regularity to leverage by a pretext task. Graphs, unlike images or text contain little redundancy or intrinsic semantics.
@@
In this paper, we propose a new model, named GOAL (for Generalist combinatorial Optimization Agent Learning), which consists of a single backbone and light-weight problem-specific adapters for input and output processing. To handle the large discrepancy between CO problems, we propose three complementary strategies. First, to ensure learning related representations for different CO problems, we propose to use a problem-specific projection of the input instance features to a small fixed-size representation, which is then projected into the main embedding space using a shared codebook. Second, to handle problems defined on graphs with arbitrary combinations of node, edge and instance-level features, we design a new form of mixed-attention blocks for the backbone, injecting edge information at the core of each attention block. Finally, we introduce a novel multi-type transformer architecture to deal with problems which involve heterogeneous types of nodes or edges (typically represented as multipartite graphs). The idea is to duplicate the backbone’s mixed-attention blocks in order to attend each meaningful combination of types while relying on the same fixed set of shared parameters. In principle, GOAL is able to handle any CO problem whose instances can be represented as a graph and where feasible solutions can be constructed iteratively, as a sequence of node selections. We find that this pattern is generic enough to cover a wide range of problems. We train GOAL in a multi-task setting, by imitation of expert trajectories of eight classic CO problems, spanning multiple domains: Routing problems: the Asymmetric Traveling Salesman Problem, the Capacitated Vehicle Routing Problem with or without Time Windows, the Orienteering Problem ; Scheduling problems: the Job Shop Scheduling Problem and the Unrelated Machine Scheduling Problem; Packing problems: the Knapsack Problem; and Graph problems: the Minimum Vertex Covering Problem. The expert trajectories are based on good-quality solutions, easily generated by traditional problem-specific solvers for instances of medium size (generally 100 nodes).
@@
Experimentally we first show that GOAL performs very well on all the training tasks, making it the first model that solves such a variety of CO problems. Second, we consider eight new problems and demonstrate that GOAL can be fine-tuned to reach a good performance, either in a few minutes using few-shot imitation learning or in a couple of hours without supervision. In both cases, the fine-tuning takes a fraction of the time and data needed to train a model from scratch. We further provide an ablation study to evaluate the effectiveness of a shared codebook, multi-type transformer architecture and mixed-attention blocks. Notably the GOAL architecture gives state-of-the-art results on 7 out of our 8 training tasks when used in as single-task model. In summary, our contributions are the following:
@@ 2 Related Works
Our work falls into the category of end-to-end constructive heuristics for combinatorial optimization. Constructive neural heuristics have been successfully applied to a wide range of COPs, including many variants of routing problems [kool_attention_2019, xin_multi-decoder_2021, kim_sym-nco_2022, kool_deep_2021, fu_generalize_2021, sun_difusco_2023], scheduling problems [zhang_learning_2020, kwon_matrix_2022], as well as classic COPs such as the Knapsack Problem [kwon_pomo_2020] and the Maximum Independent Set [khalil_learning_2017, sun_difusco_2023]. Another successful category of neural CO methods are improvement heuristics, which focus on iteratively enhancing a given complete solution, typically by using a neural network to guide the selection of local search operators [chen_learning_2019, hottung_neural_2019, ma_learning_2021, wu_learning_2020]. The seminal work of [khalil_learning_2017] introduces graph embeddings which do work for a variety of CO problems but with limited performance. Current state-of-the-art models are tailored for certain problem structures (e.g. euclidian routing problems, where nodes are represented by their coordinates plus some extra node or graph features such as demand or vehicle capacity), and successfully handled by transformer-based architectures [kool_attention_2019]. For non-euclidian routing problems, [kwon_matrix_2022] and [drakulic_bq-nco_2023] propose adapted architectures to deal with the distance matrix, but they are specialized for such problems and do not apply to those with node-only features. In GOAL, we propose a novel mixed-attention modules which can flexibly deal with both node and edge features, incurring no overhead when one or the other is missing.
@@
While the first seminal models for neural CO had a dramatic drop in performance when tested on out-of-training distribution instances [joshi_learning_2022, manchanda_generalization_2022], several recent approaches show a greatly improved generalization performance, especially to larger instances [fu_generalize_2021, li_learning_2021, son_meta-sage_2023, pan_h-tsp_2023], to distribution shifts of other instance parameters [bi_learning_2022], or both [manchanda_generalization_2022, bi_learning_2022, zhou_towards_2023, drakulic_bq-nco_2023]. Concurrently with our work, some papers have started to explore the cross-problem generalization for the family of routing problems. Specifically, [lin_cross-problem_2024] propose to pretrain the AM model [kool_attention_2019] on the TSP and to fine-tune either the entire model or some light-weight modules on the Price Collecting TSP and OP. Additionally, [zhou_mvmoe_2024, liu_multi-task_2024, berto_routefinder_2024] explore generalization to different variants of the CVRP. They build on the idea of zero-shot compositionality [ruis_independent_2021] by training a model on CVRPs with some combinations of constraints from a predefined set (e.g. time windows and duration limit constraints) and show that the model generalizes to some unseen combinations of the same constraints. Although these approaches tackle multiple problems, it is important to note that the considered Euclidian routing problems remain very similar in structure and indeed many of those problems were solved by the seminal Attention Model of [kool_attention_2019] with little adaptations of its base model. Our paper goes further and considers multi-task training and transfer learning of a new model on a variety of COPs with different structures.
@@
Multi-task (pre-)training has recently shown promising results in related fields. For example, [ibarz_generalist_2022] proposed a graph neural network processor capable of learning to execute a variety of algorithms, such as sorting and searching. The paper focuses only on polynomial problems and learns to imitate the execution of classical algorithms. In contrast, we address NP-hard problems and learn by imitation of the solution provided by (potentially exponential) specialized solvers. Recent approaches in control rely on sequence modeling of expert trajectories from multiple tasks [stooke_decoupling_2021, sun_smart_2023]. Although they treat various tasks, with potentially diverse state and action spaces, these works still consider image observations and exploit this common structure. [reed_generalist_2022] proposes GATO, a general-purpose agent obtained by training a large transformer-based model on hundreds of tasks spanning vision, language and control. It demonstrates the potential of a single network (albeit with 1.2B parameters) to effectively tackle a wide variety of tasks.
@@ 3 Multi-Task Learning for Combinatorial Optimization
Figure 1: Left: Architecture of GOAL for single-type problems; the green components are task-specific, while the others belong to the backbone. The codebook is a shared $L{\times}D$ (resp. $\bar{L}{\times}\bar{D}$) matrix applied to the node (resp. edge) representations produced by the input adapters. Right: Architecture of GOAL for multi-type problems (with here 2 node types T1,T2); “self” and “cross” are identical multi-head mixed attention blocks (same parameters per layer for all types and tasks); “FF” for the two types are also identical Feed Forward blocks per layer. For the edges, all the type combinations are shown (here T11,T12,T21,T22) but each task can define its own. The same set of parameters is used for the two architectures.
@@ 3.1 Constructive CO problem solving
We assume that a CO problem instance is given by a pair of a finite, non-empty set $X$ of feasible solutions, and a real-valued objective function $f$ whose domain contains $X$. Solving the instance consists in finding an optimal solution, i.e. an element of $\arg\min_{x\in X}f(x)$. Constructive approaches assume that $f$ is defined not only over $X$ but over a set of partial solutions, which includes $X$ and is amenable to a step-wise construction procedure where each partial solution can be obtained by a sequence of construction steps. Thus, solving the problem instance can be formulated as a sequential decision procedure which iteratively selects the construction steps so as to reach not only a feasible but an optimal solution. That in turn can be modeled as a Markov Decision Problem (MDP) associated with the problem instance. In this paper, we use the BQ-MDPs introduced by [drakulic_bq-nco_2023] which are defined for the widespread class of tail-recursive problems  [bertsekas_dynamic_vol1_2012]. In BQ-MDPs, the actions are the construction steps, typically the selection of a node, while the states are instances.
@@ 3.3 Task Representation
Hence, the GOAL model takes three inputs $t,x,\bar{x}$, where $t$ is a task identifier, and $(x,\bar{x}){\in}\boldsymbol{\Omega}_{t}$ is a parameter representing an instance of task $t$, where $x{\in}\mathbb{R}^{N{\times}F_{t}}$ (resp. $\bar{x}{\in}\mathbb{R}^{N{\times}N{\times}\bar{F}_{t}}$) represents the feature vectors of dimension $F_{t}$ of the $N$ nodes (resp. $\bar{F}_{t}$ of the $N{\times}N$ edges). While some tasks may have no edge feature ($\bar{F}_{t}{=}0$), we always assume that at least one node feature exists, which distinguishes the nodes by a randomly assigned value between 00 and $1$, acting as node identifier, so we always have $F_{t}{\geq}1$. Both $F_{t}$ and $\bar{F}_{t}$, as well as the meaning of the corresponding features, depend on task $t$. The output of the model is a matrix $y{\in}\mathbb{R}^{N{\times}K_{t}}$ representing the score of each candidate action, i.e. the pair of a node and an option for selection. As usual, it is passed to a softmax (over the whole matrix) to obtain a probability distribution. Wherever appropriate, task-dependent masking can be used to ensure that probabilities are null for the actions disallowed by a task.

### v3

# GOAL: A Generalist Combinatorial Optimization Agent Learner

@@ Abstract
Machine Learning-based heuristics have recently shown impressive performance in solving a variety of hard combinatorial optimization problems (COPs). However they generally rely on a separate neural model, specialized and trained for each single problem. Any variation of a problem requires adjustment of its model and re-training from scratch. In this paper, we propose GOAL (for Generalist combinatorial Optimization Agent Learner), a generalist model capable of efficiently solving multiple COPs and which can be fine-tuned to solve new COPs. GOAL consists of a single backbone plus light-weight problem-specific adapters for input and output processing. The backbone is based on a new form of mixed-attention blocks which allows to handle problems defined on graphs with arbitrary combinations of node, edge and instance-level features. Additionally, problems which involve heterogeneous types of nodes or edges are handled through a novel multi-type transformer architecture, where the attention blocks are duplicated to attend the meaningful combinations of types while relying on the same shared parameters. We train GOAL on a set of routing, scheduling and classic graph problems and show that it is only slightly inferior to the specialized baselines while being the first multi-task model that solves a wide range of COPs. Finally we showcase the strong transfer learning capacity of GOAL by fine-tuning it on several new problems. Our code is available at this url.
@@ 1 Introduction
Combinatorial Optimization (CO) problems are extensively studied in the Operations Research community, because they are both challenging in terms of theoretical complexity and central in many applications such as transportation, logistics and finance. Traditional CO methods include exact solvers, typically based on integer programming [korte_combinatorial_2018], and approximation algorithms [vazirani_approximation_2003] which provide optimality or approximation guarantees, respectively, but do not scale; and heuristic algorithms [boussaid_survey_2013] which produce good-quality solutions faster but without guarantees. Effective heuristics heavily rely on expert knowledge and are generally problem-specific. A more recent and promising family of methods for CO is based on Machine Learning, either to learn the hyper-parameters of classic CO solvers, or to directly learn heuristics in an end-to-end fashion [bengio_machine_2021, cappart_combinatorial_2021]. In particular constructive neural heuristics have been successful at solving a variety of CO problems, including routing [bello_neural_2017, nazari_reinforcement_2018, kool_attention_2018, kwon_pomo_2020], scheduling [zhang_learning_2020, kwon_matrix_2021] and graph problems [khalil_learning_2017, qiu_dimes_2022]. However, in most cases, a separate model is specialized and trained for each problem, and sometimes even for each distribution of instances within a problem, especially for varying instance sizes. While much effort has been devoted to improving the generalization to different instance distributions [manchanda_generalization_2022, zhou_towards_2023, son_meta-sage_2023, drakulic_bq-nco_2023], generalization to different problems is less studied. Indeed, although there are strong neural heuristics for many CO problems, any variation of these problems requires adjustment of their model and re-training from scratch. While it is true that some problem-specific model tailoring may be required, since the structure of problem instances and solutions are very diverse, many problems are related and it would be interesting to exploit their commonalities. In this paper, we seek to answer the following question: Can we design and train a single generalist model so that it can be efficiently adapted to solve a variety of CO problems? The idea of pretraining a unified model on some pretext task so that it can be efficiently fine-tuned to several downstream tasks has been extremely successful in the language and vision domains [devlin_bert_2019, brown_language_2020, radford_learning_2021]. In these domains, a self-supervised pretraining task exploits the intrinsic structure and regularities of images and text to learn generic representations. In contrast, in CO, while most problem instances can be modeled as graphs, there is no generic regularity to leverage by a pretext task. Graphs, unlike images or text contain little redundancy or intrinsic semantics.
@@
In this paper, we propose a new model, named GOAL (for Generalist combinatorial Optimization Agent Learner), which consists of a single backbone and light-weight problem-specific adapters for input and output processing. To handle the large discrepancy between CO problems, we propose three complementary strategies. First, to ensure learning related representations for different CO problems, we propose to use a problem-specific projection of the input instance features to a small fixed-size representation, which is then projected into the main embedding space using a shared codebook. Second, to handle problems defined on graphs with arbitrary combinations of node, edge and instance-level features, we design a new form of mixed-attention blocks for the backbone, injecting edge information at the core of each attention block. Finally, we introduce a novel multi-type transformer architecture to deal with problems which involve heterogeneous types of nodes or edges (typically represented as multipartite graphs). The idea is to duplicate the backbone’s mixed-attention blocks in order to attend each meaningful combination of types while relying on the same fixed set of shared parameters. In principle, GOAL is able to handle any CO problem whose instances can be represented as a graph and where feasible solutions can be constructed iteratively, as a sequence of node selections. We find that this pattern is generic enough to cover a wide range of problems. We train GOAL in a multi-task setting, by imitation of expert trajectories of eight classic CO problems, spanning multiple domains: Routing problems: the Asymmetric Traveling Salesman Problem, the Capacitated Vehicle Routing Problem with or without Time Windows, the Orienteering Problem ; Scheduling problems: the Job Shop Scheduling Problem and the Unrelated Machine Scheduling Problem; Packing problems: the Knapsack Problem; and Graph problems: the Minimum Vertex Covering Problem. The expert trajectories are based on good-quality solutions, easily generated by traditional problem-specific solvers for instances of medium size (generally 100 nodes).
@@
Experimentally we first show that GOAL performs very well on all the training tasks, making it the first model that solves such a variety of CO problems. Second, we consider eight new problems and demonstrate that GOAL can be fine-tuned to reach a good performance, either in a few minutes using few-shot imitation learning or in a couple of hours without supervision. In both cases, the fine-tuning takes a fraction of the time and data needed to train a model from scratch. We further provide an ablation study to evaluate the effectiveness of a shared codebook, multi-type transformer architecture and mixed-attention blocks. Notably the GOAL architecture gives state-of-the-art results on 7 out of our 8 training tasks when used as single-task model. In summary, our contributions are the following:
@@ 2 Related Works
Our work falls into the category of end-to-end constructive heuristics for combinatorial optimization. Constructive neural heuristics have been successfully applied to a wide range of COPs, including many variants of routing problems [kool_attention_2018, xin_multi-decoder_2021, kim_sym-nco_2022, kool_deep_2022, fu_generalize_2021, sun_difusco_2023], scheduling problems [zhang_learning_2020, kwon_matrix_2021], as well as classic COPs such as the Knapsack Problem [kwon_pomo_2020] and the Maximum Independent Set [khalil_learning_2017, sun_difusco_2023]. Another successful category of neural CO methods are improvement heuristics, which focus on iteratively enhancing a given complete solution, typically by using a neural network to guide the selection of local search operators [chen_learning_2019, hottung_neural_2020, ma_learning_2021, wu_learning_2022]. The seminal work of khalil_learning_2017 introduces graph embeddings which do work for a variety of CO problems but with limited performance. Current state-of-the-art models are tailored for certain problem structures (e.g. euclidian routing problems, where nodes are represented by their coordinates plus some extra node or graph features such as demand or vehicle capacity), and successfully handled by transformer-based architectures [kool_attention_2018]. For non-euclidian routing problems, kwon_matrix_2021 and drakulic_bq-nco_2023 propose adapted architectures to deal with the distance matrix, but they are specialized for such problems and do not apply to those with node-only features. In GOAL, we propose a novel mixed-attention modules which can flexibly deal with both node and edge features, incurring no overhead when one or the other is missing.
@@
While the first seminal models for neural CO had a dramatic drop in performance when tested on out-of-training distribution instances [joshi_learning_2022, manchanda_generalization_2022], several recent approaches show a greatly improved generalization performance, especially to larger instances [fu_generalize_2021, li_learning_2021, son_meta-sage_2023, pan_h-tsp_2023], to distribution shifts of other instance parameters [bi_learning_2022], or both [manchanda_generalization_2022, bi_learning_2022, zhou_towards_2023, drakulic_bq-nco_2023]. Concurrently with our work, some papers have started to explore cross-problem generalization for the family of routing problems. Specifically, [lin_cross-problem_2024] propose to pretrain the AM model [kool_attention_2018] on the TSP and to fine-tune either the entire model or some light-weight modules on the Price Collecting TSP and OP. Additionally, zhou_mvmoe_2024, liu_multi-task_2024, berto_routefinder_2024 explore generalization to different variants of the CVRP. They build on the idea of zero-shot compositionality [ruis_independent_2021] by training a model on CVRPs with some combinations of constraints from a predefined set (e.g. time windows and duration limit constraints) and show that the model generalizes to some unseen combinations of the same constraints. wang_efficient_2023 propose a bandit based approach to alternate between problems during training, so as to focus on those which have the best return. Although these approaches tackle multiple problems, it is important to note that the considered Euclidian routing problems remain very similar in structure and indeed many of those problems were solved by the seminal Attention Model of [kool_attention_2018] with little adaptations of its base model. Our paper goes further and considers multi-task training and transfer learning of a new model on a variety of COPs with different structures.
@@
Multi-task (pre-)training has recently shown promising results in related fields. For example, [ibarz_generalist_2022] proposed a graph neural network processor capable of learning to execute a variety of algorithms, such as sorting and searching. The paper focuses only on polynomial problems and learns to imitate the execution of classical algorithms. In contrast, we address NP-hard problems and learn by imitation of the solution provided by (potentially exponential) specialized solvers. Recent approaches in control rely on sequence modeling of expert trajectories from multiple tasks [stooke_decoupling_2021, sun_smart_2023]. Although they treat various tasks, with potentially diverse state and action spaces, these works still consider image observations and exploit this common structure. reed_generalist_2022 propose GATO, a general-purpose agent obtained by training a large transformer-based model on hundreds of tasks spanning vision, language and control. It demonstrates the potential of a single network (albeit with 1.2B parameters) to effectively tackle a wide variety of tasks.
@@ 3 Multi-Task Learning for Combinatorial Optimization
Figure 1: Left: Architecture of GOAL for single-type problems; the green components are task-specific, while the others belong to the backbone. The codebook is a shared ×$\ell{\times}D$ (resp. ¯×$\bar{\ell}{\times}\bar{D}$) matrix (with $\ell{\ll}D,\bar{\ell}{\ll}\bar{D}$) applied to the node (resp. edge) representations produced by the input adapters. Right: Architecture of GOAL for multi-type problems (with here 2 node types T1,T2); it uses multiple mixed attention blocks (“MMA”), some in self- and some in cross-attention mode, but they all share the same parameters (per layer); so do the two Feed Forward blocks (“FF”). For the edges, all the type combinations are shown (here T11,T12,T21,T22) but those which have no meaning in a task are simply omitted. The same set of parameters is used for the two architectures.
@@ 3.1 Constructive CO problem solving
We assume that a CO problem instance is given by a pair of a finite, non-empty set $X$ of feasible solutions, and a real-valued objective function $f$ whose domain contains $X$. Solving the instance consists in finding an optimal solution, i.e. an element of $\arg\min_{x\in X}f(x)$. Constructive approaches assume that $f$ is defined not only over $X$ but over a set of partial solutions, which includes $X$ and is amenable to a step-wise construction procedure where each partial solution can be obtained by a sequence of construction steps. Thus, solving the problem instance can be formulated as a sequential decision procedure which iteratively selects the construction steps so as to reach not only a feasible but an optimal solution. That in turn can be modeled as a Markov Decision Problem (MDP) associated with the problem. In this paper, we use the BQ-MDPs introduced by drakulic_bq-nco_2023 which are defined for “tail-recursive” problems, i.e. the widespread class of problems that are solvable by dynamic programming [bertsekas_reinforcement_2019]. In BQ-MDPs, the actions are the construction steps, typically the selection of a node, while the states are simply instances. This differs from many peer NCO models where each state is a pair of an instance and a partial solution. The tail recursion property precisely ensures that the two alternatives lead to “bisimilar” MDPs, which have the same optimal policies, as shown in drakulic_bq-nco_2023, but the BQ-MDPs’ states more efficiently capture the relevant information to determine these policies.
@@ 3.3 Task Representation
Hence, the GOAL model takes three inputs $t,x,\bar{x}$, where $t$ is a task identifier, and $(x,\bar{x}){\in}\boldsymbol{\Omega}_{t}$ is a parameter representing an instance of task $t$, where $x{\in}\mathbb{R}^{N{\times}F_{t}}$ (resp. $\bar{x}{\in}\mathbb{R}^{N{\times}N{\times}\bar{F}_{t}}$) represents the feature vectors of dimension $F_{t}$ of the $N$ nodes (resp. $\bar{F}_{t}$ of the $N{\times}N$ edges). While some tasks may have no edge feature ($\bar{F}_{t}{=}0$), we always assume that at least one node feature exists, which distinguishes the nodes by a randomly assigned value between 00 and $1$, similar to the random features used in Graph Neural Networks [sato_random_2021], so we always have $F_{t}{\geq}1$. Both $F_{t}$ and $\bar{F}_{t}$, as well as the meaning of the corresponding features, depend on task $t$. The output of the model is a matrix $y{\in}\mathbb{R}^{N{\times}K_{t}}$ representing the score of each candidate action, i.e. the pair of a node and an option for selection. As usual, it is passed to a softmax (over the whole matrix) to obtain a probability distribution. Wherever appropriate, task-dependent masking can be used to ensure that probabilities are null for the actions disallowed by a task.